OpenStudy (anonymous):

INTEGRATION! (Will medal) I am having some issues with the following integral.

OpenStudy (anonymous):

\[\int\limits_{?}^{?}t (2t+3)^{1/2} - 3t\]

OpenStudy (anonymous):

I know I should use the 'u method' but am having issues with the 3t. Should I solve 'u' for t and plug into 3t?

OpenStudy (kainui):

Actually the substitution part is the hardest part, all you have to do to simplify your integral is break it apart before you do that, here like this: \[\int t (2t+3)^{1/2} - 3t dt=\int t (2t+3)^{1/2} dt + \int - 3t dt\] Now you can just solve them separately :)

OpenStudy (anonymous):

I overthought that hahaha! Thank you so much!

OpenStudy (kainui):

Yeah glad to help! :D